4.5(x-2)+1.5x=2(3x-4)-1 what is the solution and how many are there?

Found 2 solutions by stanbon, unlockmath:
Answer by stanbon(75887)

4.5(x-2)+1.5x=2(3x-4)-1
4.5x - 9 + 1.5x = 6x -8 -1
------------------
6x + 9 = 6x -9
----
That is a contradiction.
Ans: No solution
===================
Cheers,
Stan H.
===================

Answer by unlockmath(1688)

Hello,
This 4.5(x-2)+1.5x=2(3x-4)-1 we can rewrite as:
4.5x-9+1.5x=6x-8-1
Rewrite:
6x-8=6x-9
Subtract 6x and add 8 to get:
0 = -1
Obviously this is incorrect so there is:
No Solution.
Make sense?
RJ
www.math-unlock.com